# Who Pays for Fairness? Rethinking Recourse under Social Burden

### v1

@@ B.2 Fairness in Prediction
Hardt et al. [23] propose a fairness-enhancing intervention that modifies the outcomes of a given predictor to satisfy a given fairness property defined by either equality of opportunity or equality of odds. In the case of binary predictors, their method flips the decisions with a given probability to satisfy the fairness criteria. The code for this method can be found in the AIF360888https://aif360.readthedocs.io/en/stable/ Python package.
We selected POSTPRO not only for its widespread use in fairness-in-prediction benchmarking, but also because it is compatible with any base classifier, enabling a fair comparison with MISOB.
B.3 Recourse Methods
We here describe the recourse methods used in our experimental evaluation (Section 5). For all the recourse methods, we consider the default hyperparameter settings (see Table B.1) and the implementation of the benchmark by Pawelczyk et al. [38].
Growing Spheres [31]
(GS) is a random search algorithm, which generates samples around the instance under consideration using growing hyperspheres until a point with the positive class label is found.
Differentiable Recourse [49]
(WT) proposed a method that generates recommendations by minimizing an objective function using gradient descent to find a nearby ”possible world” in which the individual receives a positive outcome, measuring proximity using the $\ell_{1}$ norm.
Counterfactual Conditional Heterogeneous Variational Autoencoder [37]
(CCHVAE) generates recommendations by ensuring that the latter maps the instance into a region of high data density that is as close as possible to the original instance, where proximity is computed in a lower-dimensional latent space defined by the VAE.
Table B.1: The hyperparameter values used in the different recourse methods.
@@ B.3 Recourse Methods
loss_type string ’MSE’
@@
B.4 Additional Training Details
@@
To mitigate overfitting during training, we employed mini-batches both in the regular training and iterative procedures for both NN and linear regression (LR). Specifically, we used a batch size of 256, a learning rate of 0.001, and the Adam optimizer. The base classifiers were trained for 6 epochs. In the case of POSTPRO, we performed the fairness-enhancing intervention after this training phase. For the MISOB framework, we implemented a two-stage training strategy: a warm-up phase of 3 epochs, followed by 3 additional epochs of iterative optimization. We experimented with longer training schedules but did not observe substantial performance gains. Therefore, we opted to limit the number of epochs to maintain training efficiency.
B.5 Additional Implementation Details
We have run all our experiments in a server with eight NVIDIA A5000 GPUs with an AMD EPYC 7252 CPU and 377GB of RAM. Nevertheless, our approach is neither RAM nor CPU/GPU intensive.
Appendix C Additional Results
@@
In this section, we present additional experimental results using the German Credit and Give Me Some Credit datasets (see Tables C.1 and C.2, respectively). In both cases, we consider age as the sensitive attribute. Alongside NN, we also evaluate linear regression (LR) as an additional base classifier.
@@
The results across these datasets and base classifiers reflect the same trends observed in the experiments from the main text with the Adult dataset. Specifically, MISOB consistently achieves the best worst-group performance, outperforming all other methods in terms of minimax fairness guarantees, both across the entire recourse pipeline and in the predictions of the base classifier, while incurring minimal or no loss in overall accuracy. Notably, although POSTPRO reduces TPR disparity at the base classifier level, it does so by lowering the TPR of the privileged group rather than improving that of the unprivileged group. As a result, the reduction in TPR gap comes at the cost of significantly higher burden and cost, and in some cases, reduced base accuracy.
@@
These additional results also offer further insights. MISOB remains effective even on smaller datasets. However, we observe a higher standard deviation in performance metrics across all methods, a phenomenon also noted in the fairness literature [2], which can complicate the extraction of reliable conclusions. Finally, especially in the case of the Give Me Some Credit dataset, we note that low data quality may hinder the training of strong base classifiers. Nevertheless, MISOB is still able to significantly reduce burden and cost, even when the base classifier itself performs poorly.
[…]
Base Recourse Fairness acc Burden (Eq( 3)) TPR Cost (Eq( 2)) AR
@@
[…]
Base Recourse Fairness acc Burden (Eq( 3)) TPR Cost (Eq( 2)) AR

### v3

@@ B.2 Fairness in Prediction
hardt2016equality propose a fairness-enhancing intervention that modifies the outcomes of a given predictor to satisfy a given fairness property defined by either equality of opportunity or equality of odds. In the case of binary predictors, their method flips the decisions with a given probability to satisfy the fairness criteria. The code for this method can be found in the AIF360999https://aif360.readthedocs.io/en/stable/ Python package. We selected POSTPRO not only for its widespread use in fairness-in-prediction benchmarking, but also because it is compatible with any base classifier, enabling a fair comparison with MISOB.
Table 3: Hyperparameters for each recourse method used in Section 5. For CCHVAE, we denote with #mutable_feat the number of actionable features available in each dataset. Please refer to the source code and to the implementation of pawelczyk1carla for additional details.
@@
loss_type string “MSE”
@@
B.3 Recourse Methods
Here, we describe in greater detail the recourse methods used in our experimental evaluation (Section 5). For all the recourse methods, we consider the default hyperparameter settings (see Table 3) and the implementation of pawelczyk1carla.
  • Growing Spheres (GS) (laugel2017inverse): it is a random search algorithm, which generates samples around the instance under consideration using growing hyperspheres until a point with the positive class label is found;
  • Differentiable Recourse (WT)(wachter2017counterfactual): it is a method that generates recommendations by minimizing an objective function using gradient descent to find a nearby “possible world” in which the individual receives a positive outcome, measuring proximity using the $\ell_{1}$ norm;
  • Counterfactual Conditional Heterogeneous Variational Autoencoder (CCHVAE) (pawelczyk2020learning): it generates recommendations by ensuring that the latter maps the instance into a region of high data density that is as close as possible to the original instance, where proximity is computed in a lower-dimensional latent space defined by the VAE.
B.4 Additional Training and Implementation Details
@@
We employed mini-batches both in the regular training and iterative procedures for both NN and linear regression (LR) to mitigate overfitting during training. Specifically, we used a batch size of 256, a learning rate of 0.001, and the Adam optimizer. The base classifiers were trained for 6 epochs. In the case of POSTPRO, we performed the fairness-enhancing intervention after this training phase. For the MISOB framework, we implemented a two-stage training strategy: a warm-up phase of 3 epochs, followed by 3 additional epochs of iterative optimization. We experimented with longer training schedules but did not observe substantial performance gains. Therefore, we opted to limit the number of epochs to maintain training efficiency. Lastly, we have run all our experiments in a server with eight NVIDIA A5000 GPUs with an AMD EPYC 7252 CPU and 377GB of RAM. Nevertheless, our approach is neither RAM nor CPU/GPU intensive.
Appendix C Further Results on Additional Classification Tasks
@@
Table 4 and 5 present additional experimental results using the German Credit and GiveMeSomeCredit datasets, respectively. In both cases, we consider age as the sensitive attribute. Alongside a feed-forward neural network (NN), we also evaluate a simple linear regression (LR) classifier as an additional base model.
@@
The results across these datasets and base classifiers reflect the same trends observed in the experiments from the main text with the Adult dataset. Specifically, MISOB consistently achieves the best worst-group performance, outperforming all other methods in terms of minimax fairness guarantees, both across the entire recourse pipeline and in the predictions of the base classifier, while incurring minimal or no loss in overall accuracy. Notably, although POSTPRO reduces TPR disparity at the base classifier level, it does so by lowering the TPR of the privileged group rather than improving that of the unprivileged group. As a result, the reduction in TPR gap comes at the cost of a higher burden and cost, and in some cases, reduced base accuracy.
@@
These additional results also offer further insights. MISOB remains effective even on smaller datasets. However, we observe a higher standard deviation in performance metrics across all methods, a phenomenon also noted in the fairness literature (barrainkua2024uncertainty), which can complicate the extraction of reliable conclusions. Finally, especially in the case of the GiveMeSomeCredit dataset, we note that low data quality may hinder the training of strong base classifiers. Nevertheless, MISOB is still able to significantly reduce burden and cost, even when the base classifier itself performs poorly.
Table 4: Empirical average test results on the German credit dataset across different base classifiers and recourse methods. The best results for each setting are highlighted in bold, and we report whether the target metric should increase ($\uparrow$}) or decrease ($\downarrow$}). We set age as the sensitive attribute. For MISOB, we set $\alpha=0.3$ in all experiments (cf. Eqn. 4). Each cell shows the average and standard deviation over 10 runs.
Base Recourse Fairness acc Burden (Eq. 3) TPR Cost (Eq. 2) AR
@@
Table 5: Empirical average test results on the GiveMeSomeCredit dataset across different base classifiers and recourse methods. The best results for each setting are highlighted in bold, and we report whether the target metric should increase ($\uparrow$}) or decrease ($\downarrow$}). We set age as the sensitive attribute. For MISOB, we set $\alpha=0.3$ in all experiments (cf. Eqn. 4). Each cell shows the average and standard deviation over 10 runs.
Base Recourse Fairness acc Burden (Eq. 3) TPR Cost (Eq. 2) AR